Circular No. 7013

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 F. E. Marshall, Goddard Space Flight Center; D. A. Smith, Mas-

sachusetts Institute of Technology; and T. Dotani and Y. Ueda, Institute of Space and Astronautical Science, on behalf of the ASCA team, report: “Using an ASCA observation on Sept. 12, we find an improved x-ray position for this transient (IAUC 7008) of α = 15^{h}50^{m}59^{s}, δ = −56^{o}29′.2 (equinox 2000.0) with an estimated uncertainty of 0′.6 (90-percent confidence). This position is consistent with the positions of the possible optical and radio counterparts reported on IAUC 7009 and 7010. The average 2–10-keV flux was ∼ 2.3 × 10^{−8} ergs s^{−1} cm^{−2}.”

A. J. Castro-Tirado, Instituto de Astrofísica de Andalucía, Granada, and Laboratorio de Astrofísica Espacial y Física Fundamental, Madrid; H. W. Duerbeck, University of Münster; I. Hook, European Southern Observatory (ESO), Garching; and L. Yan, Carnegie Observatories, Pasadena, report: “A 300-s optical spectrum (range 350–740 nm, resolution 2.0 nm) obtained with the 3.6-m telescope (+ EFOSC2) at ESO, La Silla, on Sept. 15.98 UT, of the proposed optical counterpart (IAUC 7009) shows a strong and broad emission line arising from Hα (656.3 nm, FWHM = 1250 km/s, equivalent width = 1.2 nm), as well as broad and weaker emission lines from Hβ (486.1 nm, FWHM = 1650 km/s, EW = 0.4 nm) and He II (468.6 nm, FWHM = 1500 km/s, EW = 0.4 nm). N III (464.0 nm) is also marginally detected. These lines are typical of soft x-ray transients in outburst. The emission lines and several interstellar absorption lines are superposed on a red continuum. From observations carried out with the Dutch 0.91-m telescope at La Silla in the Bessel V and Gunn i filters, we obtain V = 16.7 ± 0.1, V − i = +1.80 (Sept. 16.0). Additional spectroscopy and photometry during the outburst would be highly valuable.”

COMET 88P/HOWELL Visual m_{1} estimates: Aug. 10.43 UT, 11.7 (M. Mattiazzo, Wallaroo, S.

Australia, 0.20-m reflector); 22.88, 10.8 (J. Carvajal, Avila, Spain, 0.32-m reflector); Sept. 10.45, 10.8 (Mattiazzo); 13.44, 10.2 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars); 15.41, 10.4 (Seargent).

COMET P/1998 QP_{54}

Corrigendum. On IAUC 7012, line 14, for 135^{o} read 225^{o}

1998 September 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green